# Response of liquid metal in a fusion reactor blanket to rapid variation of magnetic field during a transient plasma event

### v1

@@ Abstract
Transient plasma events, such as plasma disruptions, are anticipated in the future magnetic-confinement nuclear fusion reactors. The events are accompanied by a rapid change in the magnetic field generated by the plasma current and, accordingly, induction of strong eddy currents and Lorentz forces within the reactor structure. This work targets processes within liquid-metal components of the reactor’s breeding blankets. Order-of-magnitude analysis and numerical simulations are performed to understand the response of liquid metal to a rapidly changing magnetic field and to evaluate the accuracy of commonly used simplifying model assumptions. The response is found to consist of two stages: an initial brief stage ($\sim 1$ ms) characterized by a rapid increase in the induced currents, forces, and fluid velocity; and a subsequent stage, which is triggered by the growing velocity of the metal and marked by reversals of Lorentz force, and oscillations and decreases in the amplitude of the induced fields. The transition to the second stage sets the upper limit of the velocity ($\sim 0.5$ m/s in our tests), to which an initially quiescent metal can be accelerated during the event. The simulations indicate that many widely used model assumptions, such as the negligible role of Joule dissipation in the heat balance and the constancy of physical property coefficients, remain valid during the response. However, the assumption of liquid metal incompressibility is found to be questionable due to the potential significant effects of pressure waves.
@@ I Introduction
The nature and extent of the blanket response to transient plasma events are far from clear at the moment. One major reason for this is the physical complexity of the process, which makes it challenging to model accurately. Furthermore, no experimental studies that approximate the conditions of a fusion reactor are currently feasible. Attempts of numerical simulations of an entire blanket or its module have been made (see, e.g., [4, 5]), but they inevitably use oversimplified models and suffer the effects of insufficient numerical resolution.
@@ III.2 Model
The model can be written using the magnetic induction equation (2) combined with the Ampère law (14) to calculate $\mathbf{J}$. For our simulations, we use the representation in terms of the magnetic vector potential $\mathbf{A}$, such that $\mathbf{B}=\nabla\times\mathbf{A}$ and $\mathbf{E}_{i}=-\partial\mathbf{A}/\partial t$. For convenience, the potential is separated into components corresponding to the induced and imposed nagnetic field: $\mathbf{A}=\mathbf{A}_{0}+\mathbf{a}$, where $\mathbf{B}_{0}=\nabla\times\mathbf{A}_{0}$ and $\mathbf{b}=\nabla\times\mathbf{a}$. The potential satisfies
@@
Two versions of the electromagnetic boundary conditions are used. One is the simplest available option. We solve (19) only within the liquid-metal cavity (the interior domain in figure 1) and impose the so-called pseudo-vacuum boundary conditions on the induced magnetic field
@@
where $\mathbf{n}$ is the normal to the boundary. The condition implies that the tangential component $\mathbf{b}_{\tau}$ and the normal component $J_{n}$ are zero at the boundary. Since $\nabla\cdot\mathbf{b}=0$, (20) also implies $\partial b_{n}/\partial n=0$ at the boundary.
Condition (20) is often applied in exploratory or benchmark solutions when the tangential component of the magnetic field at the boundary of a conducting region can be assumed to be zero (see, e.g., [15]) or in MHD problems at moderate magnetic Reynolds numbers, when $|\mathbf{b}|\ll|\mathbf{B}_{0}|$ (see, e.g., [16]). However, in the general case, the validity of the pseudo-vacuum condition (20) cannot be assumed a priori.
A more accurate and broadly applicable solution must include the induced magnetic field both inside and outside the conducting domain. This is done in the second version of the boundary conditions used in our study. The domain in which (19) is solved is extended on all sides by a finite non-conducting buffer zone (the exterior domain in figure 1). Conditions (20) are imposed at the outer boundary of the exterior domain. Due to the rapid decay of $\mathbf{b}$ with distance from the conducting zone, an accurate approximation by this approach requires a moderately large extension. Our test calculations show that an extension to $(L_{2},W_{2},H_{2})\geq 3(L_{1},W_{1},H_{1})$ is sufficient. A further increase in $L_{2}$, $W_{2}$, or $H_{2}$ does not affect the solution within the interior domain.
@@
where $\mathbf{F}$ is the Lorentz force (12). The other hydrodynamic effects: viscosity, advective transport, pressure, and boundary conditions on velocity at solid walls, are ignored. Such a curtailed modeling approach is justified by the focus of our study on the immediate (on the time scale $\tau\sim 10^{-2}$ s) response of the liquid to the strong and unsteady Lorentz force. The neglected hydrodynamic phenomena appear as a consequence of this response, but evolve on larger typical time scales, such as the advection $\tau_{a}\sim L/U$ and viscous $\tau_{\nu}\sim L^{2}/\nu$ scales. Another probably more important argument is that at the current level of understanding, including the full hydrodynamic behavior would have to be based on an arbitrarily chosen hydrodynamic model. As we have discussed previously, selection of the model is not self-evident and requires that the induced electric currents and the forces acting on the liquid are evaluated first.
@@ III.3 Numerical method
The finite element solver COMSOL Multiphysics® is utilized. The A-formulation of the Maxwell equations in the magnetic-field interface is adapted to match the models described in section III.2.
@@
The grid sensitivity study has determined that the interior grid with $N_{x}\times N_{y}\times N_{z}=100\times 30\times 30$ cells and $k_{rate}=6$ is sufficient for an accurate discretization.
The simulations are conducted with the time step automatically adjusted by COMSOL Multiphysics® to ensure the relative error below $10^{-6}$. Smaller time steps are utilized in the early stage of the solution, which is characterized by a particularly rapid variation of variables. The largest time step during the simulation does not exceed 0.16 ms.
@@ IV.2 Transient response at $\alpha=0$
Figure 12: Distributions of the induced magnetic field $\mathbf{b}$ at $t=0.093$ ms. Vectors are colored by the magnitude of $|\mathbf{b}|$. Results obtained for $\alpha=0$ and $\alpha=15^{\circ}$ are shown.
@@
Figure 14: Distributions of the Lorentz force $\mathbf{F}$ at $t=0.093$ ms (top row) and $t=1.54$ ms (bottom row). Vectors are colored by the magnitude of $|\mathbf{F}|$. Results obtained for $\alpha=0$ and $\alpha=15^{\circ}$ are shown. Only the half of the domain with $0\leq x\leq 0.5$ m is shown for better visibility
@@
The dominant force component $F_{x}$ accelerates the fluid in the positive $x$ direction near the wall $y=-W_{1}/2$ and in the negative $x$ direction near the wall at $y=W_{1}/2$ (see figure 17a). The velocity $u_{x}$ increases significantly, for example, to about 0.5 m/s in the areas near the sidewalls at $t\approx 1$ ms (see figires 7 and 17b). This triggers the end of the initial growth stage and the onset of the second stage of the evolution. The key factor is the contribution $\sigma u_{x}\mathbf{e}_{x}\times B_{t}\mathbf{e}_{y}=\sigma u_{x}B_{t}\mathbf{e}_{z}$ to Ohm’s law. It is opposite to the contribution $\sigma E_{i,z}$ generated directly by $dB_{p}/dt$. Recalling that the $z$ component of $\mathbf{J}$ is the main contributor to the Lorentz force, we see how $\mathbf{F}$ is affected by $\mathbf{u}$. The new contribution $-\sigma u_{x}B_{t}^{2}\mathbf{e}_{x}$ decelerates the fluid particles that are set in motion during the initial stage. At the same time, the accelerating force $-\sigma E_{i,z}B_{t}\mathbf{e}_{x}$ decreases as $E_{i,z}$ decreases following the decay of $dB_{p}/dt$. At some moment, the entire Lorentz force component $F_{x}$ starts to act against the motion of fluid particles. A proof that this actually happens is provided by the data in figure 9. The work rate of the Lorentz force $Q_{L}$ becomes negative at $t\approx 1.1$ ms.
The solution fields during the second stage are illustrated by the snapshots of $\mathbf{F}$ and $\mathbf{u}$ at $t=1.54$ ms (see figires 14c, 15c, 16c and 17c). The transformation is more complex than described above due to the three-dimensionality effects, Lorentz forces resulting from the interaction of currents with $B_{p}$, and the effect of the ends of the domain. Nontheless, the computed fields are fully consistent with the outlined scenario of the fluid velocity causing an opposing Lorentz force and a reversal of the initial trend of the system evolution. We see in figure 7 that the amplitudes of point values of $F_{x}$ and, with a delay, of $u_{x}$ start to decrease with time. The maxima of all components of $\mathbf{J}$, $\mathbf{b}$, and $\mathbf{F}$ also begin to decay (see figure 10). A clear illustration of the transformation leading to this behavior is provided by the two-dimensional distributions of the components of $\mathbf{F}$ in figures 15a-c (the middle cross-section $x=0.5$ m) and figures 16a-c (the cross-section $x=0.01$ m near the end). We see the initial growth stage at $t=0.093$ ms in figures 15a and 16a. The transitional stage, during which $F_{x}$, while not yet reversed, is much weaker, is illustrated by the distributions at $t=1.04$ ms in figures 15b and 16b. The fully developed secondary stage, during which $F_{x}$ is reversed, is shown for $t=1.54$ ms in figures 15c and 16c.
@@ Acknowledgements.
The work of Oleg Zikanov was supported by the U.S. Department of Energy through its SciDAC program under the contract DE-AC05-00OR22725.

### v3

# Exploratory study of liquid metal response to rapid variation of applied magnetic field

@@ Abstract
Transient plasma events, such as plasma disruptions, are anticipated in the future magnetic-confinement nuclear fusion reactors. The events are accompanied by a rapid change in the magnetic field generated by the plasma current and, accordingly, induction of strong eddy currents and Lorentz forces within the reactor structure. This work targets processes within liquid-metal components of the reactor’s breeding blankets. Order-of-magnitude analysis and exploratory numerical simulations are performed to understand the response of liquid metal to a rapidly changing magnetic field and to evaluate the accuracy of commonly used simplifying model assumptions. The response is found to consist of two stages: an initial brief stage ($\sim 1$ ms) characterized by a rapid increase in the induced currents, forces, and fluid velocity; and a subsequent stage, which is triggered by the growing velocity of the metal and marked by reversals of Lorentz force, and oscillations and decreases in the amplitude of the induced fields. The transition to the second stage sets the upper limit of the velocity ($\sim 0.5$ m/s in our tests), to which an initially quiescent metal can be accelerated during the event. The simulations indicate that many widely used model assumptions, such as the negligible role of Joule dissipation in the heat balance and the constancy of physical property coefficients, remain valid during the response. However, the assumption of liquid metal incompressibility is found to be questionable due to the potential significant effects of pressure waves.
@@ I Introduction
The response of liquid metal to a rapidly changing magnetic field at the parameters relevant to the blanket environment was a subject of several earlier works. Attempts of numerical simulations of an entire blanket or its module have been made (see, e.g., [4, 5]), but they inevitably use oversimplified models and suffer the effects of insufficient numerical resolution. Some insight into the physics of the response was provided in the studies of the geometries of an infinite channel [6] and infinite duct [7]. In both cases, the liquid metal was assumed to be an incompressible fluid. The imposed magnetic field had two components, a constant transverse component and a longitudinal component decaying at a constant rate. Steady-state one-dimensional (in [6]) and steady-state and transient two-dimensional (in [7]) solutions were found. As will become clear in the following, these model settings are analogous to ours, except that in our case the duct has a finite length, solutions are three-dimensional and inherently unsteady, and a more realistic time variation of the longitudinal field is assumed. We will see that there are certain qualitative similarities between our results and the predictions of [6, 7].
The nature and extent of the blanket response to transient plasma events are far from clear at the moment. One major reason for this is the physical complexity of the process, which makes it challenging to model accurately. Furthermore, no experimental studies that approximate the conditions of a fusion reactor are currently feasible.
@@ III.2 Model
The entire set of the MHD equations (14)-(18) is shown here in order to facilitate the discussion. The actual simulations use the model based on the representation in terms of the magnetic vector potential $\mathbf{A}$, such that $\mathbf{B}=\nabla\times\mathbf{A}$ and $\mathbf{E}_{i}=-\partial\mathbf{A}/\partial t$. For convenience, the potential is separated into components corresponding to the induced and imposed magnetic field: $\mathbf{A}=\mathbf{A}_{0}+\mathbf{a}$, where $\mathbf{B}_{0}=\nabla\times\mathbf{A}_{0}$ and $\mathbf{b}=\nabla\times\mathbf{a}$. The potential satisfies
@@
The electric currents $\mathbf{J}$ are calculated using Ohm’s law (18).
Two versions of the electromagnetic boundary conditions are used. One is the simplest option available. We solve (19) only within the liquid-metal cavity (the interior domain in figure 1) and impose the so-called pseudo-vacuum boundary condition on the induced magnetic field
@@
where $\mathbf{n}$ is the normal to the boundary. The pseudo-vacuum boundary condition reduces to a Dirichlet condition when expressed via magnetic flux density, but becomes a Neumann-type condition when formulated for the magnetic vector potential. The condition implies that the tangential component $\mathbf{b}_{\tau}$ and the normal component $J_{n}$ are zero at the boundary. Since $\nabla\cdot\mathbf{b}=0$, (20) also implies $\partial b_{n}/\partial n=0$ at the boundary.
Condition (20) is often applied in exploratory or benchmark solutions when the tangential component of the magnetic field at the boundary of a conducting region can be assumed to be zero (see, e.g., [17]) or in MHD problems at moderate magnetic Reynolds numbers, when $|\mathbf{b}|\ll|\mathbf{B}_{0}|$ (see, e.g., [18]). However, in the general case, the validity of the pseudo-vacuum condition (20) cannot be assumed a priori. A more detailed discussion can be found, e.g., in [18, 19, 6].
A more accurate and broadly applicable solution must include the induced magnetic field both inside and outside the conducting domain. This is done in the second version of the boundary conditions used in our study. The domain in which (19) is solved is extended on all sides by a finite non-conducting buffer zone (the exterior domain in figure 1). Conditions (20) are imposed at the outer boundary of the exterior domain. Due to the rapid decay of $\mathbf{b}$ with distance from the conducting zone, an accurate approximation by this approach requires a moderately large extension. Similar domain extension techniques was employed in [19]. A slightly different version with the true far-field condition $\bm{b}=0$ imposed at the outer boundary of the extended domain was used in [20, 6, 18]. Our test calculations show that an extension to $(L_{2},W_{2},H_{2})\geq 3(L_{1},W_{1},H_{1})$ is sufficient. A further increase in $L_{2}$, $W_{2}$, or $H_{2}$ does not affect the solution within the interior domain.
@@
where $\mathbf{F}$ is the Lorentz force (12). The other hydrodynamic effects: mass conservation of the fluid, viscosity, advective transport, pressure force, and boundary conditions on velocity at solid walls, are ignored. Such a curtailed modeling approach is justified by the focus of our study on the immediate (on the time scale $\tau\sim 10^{-2}$ s) response of the liquid to the strong and unsteady Lorentz force. The neglected hydrodynamic phenomena appear as a consequence of this response, but evolve on larger typical time scales, such as the advection $\tau_{a}\sim L/U$ and viscous $\tau_{\nu}\sim L^{2}/\nu$ scales. Another probably more important argument is that at the current level of understanding, including the full hydrodynamic behavior would have to be based on an arbitrarily chosen hydrodynamic model. As we have discussed previously, selection of the model is not self-evident and requires that the induced electric currents and the forces acting on the liquid are evaluated first.
@@ III.3 Numerical method
The finite element solver COMSOL Multiphysics® is utilized [21]. The A-formulation of the Maxwell equations in the magnetic-field interface is adapted to match the models described in section III.2.
@@
Spatial discretization is based on linear shape functions. The grid sensitivity study has determined that the interior grid with $N_{x}\times N_{y}\times N_{z}=100\times 30\times 30$ cells and $k_{rate}=6$ is sufficient for accurate discretization.
The time discretization is by the implicit backward differentiation formula with variable discretization order and time step. The discretization order varies between 2 and 5 according to the adaptive algorithm of COMSOL Multiphysics® [21]. The upper limit on the allowed time steps is set manually by the formula for time layers $t_{i}$
$t_{i}=\left[\frac{\tanh\left(D\left(\frac{i-1}{N_{t}-1}-1\right)\right)}{\tanh(D)}+1\right]t_{final},\>i=1,\ldots,N_{t},$ (35)
where $D=1.5$ and $N_{t}=300$. According to (35), smaller time steps are utilized in the early stage of the solution, which is characterized by a particularly rapid variation of variables. The limits on the smallest and the largest time steps are 0.0453 ms and 0.2494 ms, respectively. Further refinement of the time steps is provided by the time-step adjustment tool of COMSOL Multiphysics® [21, 22]. The tool ensures that the estimated local truncation error does not exceed the relative tolerance of $10^{-6}$.
The time layers (35) are also used to save the solution variables for further analysis.
@@ IV.2 Transient response at $\alpha=0$
Figure 12: Distributions of the induced magnetic field $\mathbf{b}$ at $t=0.093$ ms. Vectors are colored and scaled by the magnitude of $|\mathbf{b}|$. Magnetic flux density displayed on a uniformly spaced virtual grid comprising 1500 sample points. Results obtained for $\alpha=0$ and $\alpha=15^{\circ}$ are shown.
@@
Figure 14: Distributions of the Lorentz force $\mathbf{F}$ at $t=0.093$ ms (top row) and $t=1.54$ ms (bottom row). Vectors are colored by the magnitude of $|\mathbf{F}|$. The distributions are shown in five cross-sections $x=const$. Semi-transparent surfaces wrapping the vector and colored by the magnitude of $|\mathbf{F}|$ are shown for better visibility. Results obtained for $\alpha=0$ and $\alpha=15^{\circ}$ are shown. Only the half of the domain with $0\leq x\leq 0.5$ m is depicted.
@@
The dominant force component $F_{x}$ accelerates the fluid in the positive $x$ direction near the wall $y=-W_{1}/2$ and in the negative $x$ direction near the wall at $y=W_{1}/2$ (see figure 17a). The velocity $u_{x}$ increases significantly, for example, to about 0.5 m/s in the areas near the sidewalls at $t\approx 1$ ms (see figures 7 and 17b). This triggers the end of the initial growth stage and the onset of the second stage of the evolution. The key factor is the contribution $\sigma u_{x}\mathbf{e}_{x}\times B_{t}\mathbf{e}_{y}=\sigma u_{x}B_{t}\mathbf{e}_{z}$ to Ohm’s law. It is opposite to the contribution $\sigma E_{i,z}$ generated directly by $dB_{p}/dt$. Recalling that the $z$ component of $\mathbf{J}$ is the main contributor to the Lorentz force, we see how $\mathbf{F}$ is affected by $\mathbf{u}$. The new contribution $-\sigma u_{x}B_{t}^{2}\mathbf{e}_{x}$ decelerates the fluid particles that are set in motion during the initial stage. At the same time, the accelerating force $-\sigma E_{i,z}B_{t}\mathbf{e}_{x}$ decreases as $E_{i,z}$ decreases following the decay of $dB_{p}/dt$. At some moment, the entire Lorentz force component $F_{x}$ starts to act against the motion of fluid particles. A proof that this actually happens is provided by the data in figure 9. The work rate of the Lorentz force $Q_{L}$ becomes negative at $t\approx 1.1$ ms.
The solution fields during the second stage are illustrated by the snapshots of $\mathbf{F}$ and $\mathbf{u}$ at $t=1.54$ ms (see figures 14c, 15c, 16c and 17c). The transformation is more complex than described above due to the three-dimensionality effects, Lorentz forces resulting from the interaction of currents with $B_{p}$, and the effect of the ends of the domain. Nontheless, the computed fields are fully consistent with the outlined scenario of the fluid velocity causing an opposing Lorentz force and a reversal of the initial trend of the system evolution. We see in figure 7 that the amplitudes of point values of $F_{x}$ and, with a delay, of $u_{x}$ start to decrease with time. The maxima of all components of $\mathbf{J}$, $\mathbf{b}$, and $\mathbf{F}$ also begin to decay (see figure 10). A clear illustration of the transformation leading to this behavior is provided by the two-dimensional distributions of the components of $\mathbf{F}$ in figures 15a-c (the middle cross-section $x=0.5$ m) and figures 16a-c (the cross-section $x=0.01$ m near the end). We see the initial growth stage at $t=0.093$ ms in figures 15a and 16a. The transitional stage, during which $F_{x}$, while not yet reversed, is much weaker, is illustrated by the distributions at $t=1.04$ ms in figures 15b and 16b. The fully developed secondary stage, during which $F_{x}$ is reversed, is shown for $t=1.54$ ms in figures 15c and 16c.
@@
The scenario of the liquid metal response reported in this section can be compared with the $x$-independent solutions for infinite channel and duct at $dB_{p}/dt=const<0$ obtained in [6, 7]. The solutions include the full hydrodynamics of a viscous incompressible fluid, but limit the possible flow states to single-component fields $F_{x}$, $u_{x}$, and $b_{x}$, as well as the electric currents in the $y-z$ plane. Asymptotic steady-state and unsteady solutions evolving to steady states are obtained.
Despite the profoundly different problem statement, our solution shows a certain agreement with the results of [6, 7]. Considering only the first stage of the response and ignoring the effects of the ends of the domain, we find in figures 12, 14, 15, and 17 the distributions of $\bm{b}$, $\bm{F}$, and $\bm{u}$, which are qualitatively similar to the distributions predicted in [6, 7]. In particular, the dominant components of the fields are the $x$-components, and these components vary slowly along $x$. This provides a degree of validation for the idealized models of [6, 7]. There are also similarities in spatial distributions of $b_{x}$, $F_{x}$, and $u_{x}$, including the concentration of $b_{x}$ in the core of the fluid domain (see figure 12) or the fluid accelerated in the opposite direction near the opposite walls (see figure 17).
At the same time, there are major differences between our solution and the solutions of [6, 7]. One is the loss of three-dimensionality and end effects in [6, 7]. Another, probably more important difference is that the compex transient dynamics developing at a rapid and time-dependent variation of $B_{p}$ is reproduced by our model, but cannot be reproduced by the models of [6, 7]. A complete analysis of the process obviously requires a model that, at the same time, includes the full hydrodynamics and considers the three-dimensional unstable dynamics of the system.
@@ Acknowledgements.
The work of Oleg Zikanov was done under the subcontract CW52278 between the University of Michigan - Dearborn and the Oak Ridge National Laboratory as a part of the SciDAC project ‘‘Center for Simulation of Plasma - Liquid Metal Interactions in Plasma Facing Components and Breeding Blankets of a Fusion Power Reactor’’ supported by the U.S. Department of Energy under the contract DE-AC05-00OR22725.